Reports 1-1 of 1 Clear search Modify search
AdV-ISC (Commissioning up to first full interferometer lock)
casanueva, pinto, spinicelli, derossi, gosselin - 22:48 Friday 17 November 2023 (62486) Print this report
Green adjustment and CITF lock

At the beginning of the shift we took a look to the green lock trying to find the noise source, focusing in particula ron the north. We checked the demodulation phase, which seemed good. We tried to measure the OLTF of the north arm, but we couldn't inject much noise because the lock was already too noisy. For this reason we tried to increase the loop gain, which didn't have a big impact. Finally we tried to adjust the modulation amplitude, and its increase was enough to improve the noise and bring it back to the old level, see Figure 1. We saved the new values on the config file.

After this, we tried to lock the CITF and waited there for the thermal transient to finish. At that point we checked the demodulation phases of the CITF (mostly changes on the 56 by around 0.2-0.3rad but not on the 6). We had to adjust quite a lot the gains in some spots. I temporarily added a prcl gain change during CARM2MC because PRCL was reaching 50 there. Only by tuning gains we managed to reach CARM NULL 3f which was not very stable. We checked the SSFS phase, which didn't change, but we unlocked going to CARM NULL 1f (I forgot to check the phase of the 56MHz for MICH and SRCL).

After this ulock we couldn't come back to CARM NULL 3f, and we kept unlocking around CARM_MC_IR and DARM_IR. We leave the arms locked during the night

Images attached to this report
Search Help